OpenStudy (i_love_my_nieces):

Yvonne put $4,000 in a savings account. At the end of 3 years, the account had earned $960 in simple interest. a) How much does she have in her account at the en of 3 years? b)t what annual simple interest did the account grow? Show your work. c) How many more dollars would she have in her account if the interest rater were 1% greater? Show your work.

OpenStudy (anonymous):

I think it's c but I'm not 100% sure

OpenStudy (anonymous):

multiply the following: \[4,000\times 3 years\] then I have no idea what to do next

OpenStudy (i_love_my_nieces):

It's not MC okay

OpenStudy (anonymous):

MC? @I_love_my_nieces

OpenStudy (i_love_my_nieces):

Multiple choice.

OpenStudy (anonymous):

Oh I'm so sorry

OpenStudy (i_love_my_nieces):

It's fine!

OpenStudy (jdoe0001):

\(\bf i = prt\qquad i=960\qquad p=4,000\qquad t=3 \\ \quad \\ 960=4,000\cdot r\cdot 3\) so how much do you think she has in her account after t = 3, 3 years ? keep in mind that she put originally, as principal, 4,000

OpenStudy (i_love_my_nieces):

i'm gonna go eat now so if you can hold on!

OpenStudy (anonymous):

Haha

OpenStudy (jdoe0001):

make sure you eat your veggies

OpenStudy (anonymous):

Bahaha @jdoe0001

OpenStudy (i_love_my_nieces):

I did @jdoe0001